OpenStudy (anonymous):

I'm looking for some clues on what Series Convergence Test to use for this series, so far I have eliminated the Integral Test and the Limit Comparison test.

OpenStudy (anonymous):

\[\sum_{n =2}^{\infty} \frac{ 1 }{ \sqrt(n) \ln n}\]

OpenStudy (anonymous):

Have you tried the ratio test? Just a thought.

OpenStudy (anonymous):

I will try that quickly.

OpenStudy (anonymous):

Never mind, the ratio test fails.

OpenStudy (anonymous):

Could you help me try the Comparison test?

OpenStudy (anonymous):

Yes, I was just about to suggest that. Try comparing it to \[\sum_{n=2}^\infty\frac{1}{\sqrt n}\]

OpenStudy (anonymous):

That's actually the algebraic result i got for the Limit Comparison Test but ill try the comparison test

OpenStudy (anonymous):

Does that count as a p-series?

OpenStudy (anonymous):

Yes it does. p = 1/2, in the case of the series I suggested, so it diverges.

OpenStudy (anonymous):

You have to show that \[\sum_{n=2}^\infty\frac{1}{\sqrt n}<\sum_{n=2}^\infty\frac{1}{\sqrt n \ln(n)}\]

OpenStudy (anonymous):

That does not appear to be true

OpenStudy (anonymous):

Oh, my mistake. I had the inequality wrong in my mind. I'll try to see what else can be done.

OpenStudy (anonymous):

I might be wrong as well, I just took an arbitrary number and plugged it into both and it looks like 1/sqrt(n) comes out greater

OpenStudy (anonymous):

no way this converges, by eyeball test

OpenStudy (anonymous):

but i suppose that is not much help maybe you can use \(\ln(n)<\sqrt{n}\) and so \[\frac{1}{\sqrt{n}\ln(n)}>\frac{1}{n}\] so diverges via comparison test